SourMunchkin7806:

Find the inverse of the function. f(x) = 6x3 - 8 f-1(x) = cube root of x divided by six plus eight f-1(x) = cube root of quantity x minus eight divided by six border= Not a one-to-one function f-1(x) = cube root of quantity x plus eight divided by six

SourMunchkin7806:

@JustSaiyan

JustSaiyan:

I suck at math. I think Vocaloid will be on soon, though. If my stalker abilities are right, she should be on just before Ultri.

SourMunchkin7806:

Ok no prob man thanks anyways!

EvieSwan2405:

Here this should help you learn on how to do the problem, but then also Mathway is a good helping answer math problems! https://www.mathway.com/popular-problems/Precalculus/422679

SourMunchkin7806:

f of x equals four divided by x. and g of x equals four divided by x

dude:

Is your equation \(f(x) = 6x^3 - 8\)?

SourMunchkin7806:

@dude can you help with one more

dude:

That was not the answer, I was asking if that was the equation in the question

SourMunchkin7806:

in the question its f(x) equals four divided by x. and g(x) equals four divided by x

SourMunchkin7806:

the question is Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x

dude:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @SourMunchkin7806 in the question its f(x) equals four divided by x. and g(x) equals four divided by x \(\color{#0cbb34}{\text{End of Quote}}\) \(f(x)=\frac4x\\ g(x)=\frac4x\)? Your question is confusing

SourMunchkin7806:

yes tahts the equation

SourMunchkin7806:

so the questions is Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x

SourMunchkin7806:

thats all it gives me

dude:

But f(x) and g(x) are the same equation, that is throwing me off

SourMunchkin7806:

i know i dont have anything else to give im confused too

SourMunchkin7806:

@Vocaloid

dude:

\(f(g(x)) = x ~~~~\color{green}{f(x)=\frac4x~~~~g(x)=\frac4x}\) \(\large{f(x)=\frac{4}{\boxed{\frac4x}}}\)

SourMunchkin7806:

so what would it be?

dude:

\[\frac{a}{\frac{b}{c}}=\frac{a\cdot \:c}{b}\\ =\frac{4x}{4} =x\]

SourMunchkin7806:

so yes they are inverses?